An investment of $ 18,000 is growing at 5% compunded quarterly



a) calculate the accumulated amount of this investment at the end of year 1 Round of the nearest cent



b) if the intrest rate changed to 3% compounded monthly of the year 1, calculated the accumulated amount of this investment at the end of year 2 round of the nearest cent



c) Calculate the total amount of interest earned from the investment during the 2 year- period round of the nearest cent.

1
Expert's answer
2022-03-07T17:03:03-0500

"FV=PV(1+\frac{r}{n})^{n×t}"

compounded quarterly n =4,t=1


"=18000(1+\frac{0.05}{4})^{4×1}"


"=18000(1.0125)^4=\$18,917"


b)

compounded monthly n=12,t=2

"=18917(1+\frac{0.03}{12})^{12×2}"


"=18917(1.0025)^{24}=\$20,085"


c)interest earned

"I=PRT"

"=20085×0.03×2=\$1205.1"